Circular No. 7775

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 X2 J. V. Scotti, Lunar and Planetary Laboratory, reports his discovery of

a comet on CCD images taken with the 0.9-m Spacewatch telescope, noting a coma diameter of 5′′ and a tail 0′.40 long in p.a. 283^{o} (and m_{2} = 22.7). Confirming observations were made at Kleť by J. Tichá, M. Tichý, and P. Jelínek, who reported a 8′′ coma and a faint westward tail.

2001 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 14.43924 8^{h}13^{m}21.^{s}88 +16^{o}34′52′′.1  Scotti
14.46482 8 13 21.44 +16 34 51.5 19.3  _{”}
14.48687 8 13 21.08 +16 34 51.7 19.5  _{”}
14.97498 8 13 13.44 +16 34 51.7 18.4 Tichá
14.98387 8 13 13.38 +16 34 53.0  _{”}
14.98728 8 13 13.17 +16 34 51.8 18.6  _{”}

PSR J0205+6449 V. Malofeev, O. Malov, and A. Glushak, Pushchino Radio Astron-

omy Observatory, Astro Space Center, Lebedev Physical Institute, report the detection of radio emission from PSR J0205+6449 with the same period (65.68 ms) as the x-ray pulsar, which was discovered in the center of the young supernova remnant 3C 58 in August (Murray et al. 2001,

http://xxx.lanl.gov/abs/astro-ph/?0108489): “We observed during

2001 Sept.–Nov. with the Large Phased Array, a sensitive transit antenna consisting of 16 384 dipoles and an operating frequency of 110.5 ± 1.5 MHz. The mean flux density was 30 mJy, and the width of the integrated profile at 50 percent of maximum intensity is ∼ 8 ms (S/N = 19, with an integration of 37 785 pulses at 110.5 MHz). At about one half the pulsar period, there is an interpulse with 25 percent of the main pulse intensity. The estimate for the dispersion measure of 24 pc/cm^{3} places the pulsar at a distance ∼ 1100 pc, according to the Taylor and Cordes (1993, Ap.J. 411, 674) model of the interstellar medium. Confirming observations at 88 MHz were made with the east-west arm of the Wide-Band Cross-type Radiotelescope DKR-1000 (area ∼ 40 000 m^{2}) during Oct. 3, 7, 10, and 11. The width of the integrated profile is ∼ 12 ms and the flux density < 150 mJy, and we estimate that the spectral index is between 2 and 5. We derive a period of 0.0656789(1) s and a period derivative of 1.9(2) × 10^{−13} on JD 2451901.330 (in good agreement with the x-ray data of Murray et al. for the same epoch).”

2001 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green